# Can We Validate Counterfactual Estimations in the Presence of Unknown Network Interference?

### v1

@@ 3.4.1 An Illustrative Example
Consider a special case of outcome specification (3) where functions $h_{t}{}$ are equal to zero, and functions $g_{t}$ satisfy,
$g_{t}\big{(}Y^{i}_{t}{}(\bm{W}_{t}{}),W^{i}_{t+1},{\vec{X}}^{i}\big{)}=\delta+\theta Y^{i}_{t}{}(\bm{W}_{t}{})+\tau W^{i}_{t+1}+\lambda Y^{i}_{t}{}(\bm{W}_{t}{})W^{i}_{t+1}\,.$ (5)
In this case, if we denote the outcomes sample mean at time $t$ by $\nu_{t}{}$, the state evolution implies that as $N\rightarrow\infty$,
@@
where $f$ has the form $f(a,b)=\delta+\theta a+\tau b+\lambda ab$, and plase the role of the aforementioned time-invariant function. Subsequently, given observations $\bm{Y}{}(\bm{w}{})$ and $\bm{w}{}$, Algorithm 1 enables consistent estimation of CFE for any target treatment allocation $\bm{w}{}^{\prime}$, provided that $\bm{w}{}$ and $\bm{w}{}^{\prime}$ share identical columns for treatment assignment at time $t=0$ and the design set $\{p_{0},\ldots,p_{T}\}$ contains at least two distinct values [^7] A detailed analysis of a more general setting appears in Appendix 10.1..
@@
$(\hat{\delta},\hat{\theta},\hat{\tau},\hat{\lambda})$leftarrow Estimation of $(\delta,\theta,\tau,\lambda)$ using OLS in $\hat{\nu}_{t+1}{}=\delta+\theta\hat{\nu}_{t}{}+\tau\hat{p}_{t+1}{}+\lambda\hat{\nu}_{t}{}\hat{p}_{t+1}{}$
@@
3.4.2 General Estimation Theory
The above example demonstrates that our counterfactual estimation problem reduces to the consistent estimation of $f$. This relationship can be formalized for our general outcome specification (3). We provide an informal version of this result below, while the formal statement and complete proof appear in Theorem 10.1 in Appendix 10.
Theorem 3.6 (Consistency - informal statement)
Under certain regularity conditions, given consistent estimation of the mappings $f_{t}$ in the SE equation (4), any desired counterfactual evolution $\left\{\text{CFE}_{t}{(\bm{w}{}^{\prime})}\right\}_{t=0}^{T}$ can be consistently estimated.
3.4.3 Practical limitations
The effectiveness of approaches like Algorithm 1 in identifying treatment effects across various settings has been established by Shirani and Bayati, (2024). However, two significant constraints require consideration. First, the estimation procedure in the second step of Algorithm 1 faces sample size limitations due to the experiment horizon. Second, as demonstrated by Bayati et al., (2024), certain settings require accounting for more complex structures, necessitating a richer class of functions $f$. To capture these more complex settings, one approach is to use higher moments of the outcome and treatment distributions in state evolution (4).
For example, if we denote the variance of outcomes at time $t$ by $\rho_{t}{}$, and if functions $g$ and $h{}$ contain second-order terms, the state evolution becomes:
$\Big{(}\nu_{t+1}{},(\rho_{t+1}{})^{2}\Big{)}=f\Big{(}\nu_{t}{},(\rho_{t}{})^{2},p_{t+1},p_{t+1}^{2}\Big{)}\,.$ (7)
An alternative approach to capturing complex structure is to consider a richer family of functions $f$ during the estimation phase.
These considerations raise two key questions: First, how can we expand the sample size to improve estimation accuracy? Second, how can we determine the appropriate specification for the underlying experiment, similar to model selection in supervised learning? The following two sections address these questions systematically.
@@ 4 Distribution-preserving Network Bootstrap
To enable the use of more powerful supervised learning techniques, we introduce a theoretically supported resampling methodology that generates additional samples from the state evolution equation.
The methodology centers on batching experimental units into subpopulations, denoted by $\mathcal{S}$. Specifically, $\mathcal{S}$ represents a subset of units from $\{1,\ldots,N\}$ where membership is determined exclusively by the treatment allocation $\bm{W}{}$. Then, under the randomized treatment assumption—which specifies that $\bm{W}{}$ is random and independent of all other variables—each $\mathcal{S}$ constitutes a random sample from the experimental population. In this section, we demonstrate that each distinct selection of $\mathcal{S}$ yields a new observation of the state evolution equation, provided the empirical treatment distributions exhibit sufficient variation across subpopulations.
Our analysis proceeds in two stages. First, we extend the state evolution concept to derive subpopulation-specific state evolutions. We establish that our batching technique generates statistically efficient samples of the experimental state evolution while preserving its fundamental structure—hence the term distribution-preserving network bootstrap (DPNB). Second, we derive a decomposition rule that characterizes the finite-sample behavior of potential outcomes. This decomposition enables the identification of endogenous noise within the experimental data and elucidates how DPNB addresses this challenge.
@@
The theoretical frameworks presented here extend, with minimal or no modification, to arbitrary subpopulations. While our analysis focuses specifically on treatment-allocation-based subpopulations in this section, the appendices provide complete theoretical statements applicable to general subpopulation structures.
@@ 4.1 Subpopulation-specific State Evolution
We begin by extending the state evolution analysis to subpopulations of experimental units.
Theorem 4.2 (informal statement)
@@
Under the conditions of Theorem 3.5, consider a subpopulation $\mathcal{S}$ whose size $\left|{\mathcal{S}}\right|$ grows to infinity as $N\to\infty$. There exist mappings $\varphi_{t},\;t=0,\ldots,T-1$, such that:
$\mathcal{L}(Y^{\mathcal{S}}_{t+1}{})=\varphi_{t}\left(\mathcal{L}(Y_{t}{}),\mathcal{L}(W_{t+1}{}),\mathcal{L}(Y^{\mathcal{S}}_{t}{}),\mathcal{L}(W^{\mathcal{S}}_{t+1}{})\right),$ (8)
where $Y^{\mathcal{S}}_{t}{}$ and $W^{\mathcal{S}}_{t+1}{}$ represent the weak limits of outcomes and treatments for units in subpopulation $\mathcal{S}$, while $Y_{t}{}$ and $W_{t+1}{}$ are their corresponding population-level analogs as defined in Theorem 3.5.
@@
According to Equation (8), each subpopulation’s outcome distribution follows a distinct state evolution equation governed by mappings $\varphi_{t},\;t=0,\ldots,T-1$. The outcomes for units within $\mathcal{S}$ evolve through a dynamic interplay between population-level and subpopulation-specific outcomes and treatments. The formal statement of Theorem 4.2 appears in Theorem 9.8 of Appendix 9.4, where (27) explicitly defines $\varphi_{t}$ through mappings $g_{t}$ and $h_{t}{}$.
@@
To demonstrate the efficacy of DPNB, consider two subpopulations $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ such that $W^{\mathcal{S}_{1}}_{t+1}{}\neq W^{\mathcal{S}_{2}}_{t+1}{}$ in (8). Then, two sequences $Y^{\mathcal{S}_{1}}_{0}{},\ldots,Y^{\mathcal{S}_{1}}_{T}{}$ and $Y^{\mathcal{S}_{2}}_{0}{},\ldots,Y^{\mathcal{S}_{2}}_{T}{}$ yield distinct observations of the state evolution described by (8). Furthermore, for each time step $t$, the mapping $\varphi_{t}$ exhibits fundamental similarities to $f_{t}$, with the experimental state evolution equation in (4) emerging as a special case of (8) when applied to the entire population. Accordingly, by strategically selecting subpopulations with distinct treatment allocations, we obtain multiple state evolution observations, enabling the estimation of $\varphi_{t}$ and, consequently, $f_{t}$.
@@
Consider two distinct subpopulations: treated units ($\mathcal{S}={\cal T}$) and control units ($\mathcal{S}={\cal C}$). From (8), we have:
$\mathcal{L}(Y^{{\cal T}}_{t+1}{})=\varphi_{t}\left(\mathcal{L}(Y_{t}{}),\mathcal{L}(W_{t+1}{}),\mathcal{L}(Y^{{\cal T}}_{t}{}),1\right),$ (9)
$\mathcal{L}(Y^{{\cal C}}_{t+1}{})=\varphi_{t}\left(\mathcal{L}(Y_{t}{}),\mathcal{L}(W_{t+1}{}),\mathcal{L}(Y^{{\cal C}}_{t}{}),0\right).$
When the direct treatment effect is non-zero, the outcome distributions for the treated group $Y^{{\cal T}}_{t}{}$ and the control group $Y^{{\cal C}}_{t}{}$ are distinct, providing two different observations from the state evolution.
4.2 Outcomes Decomposition Rule
@@
Suppose Assumption 3.3 holds and let $\vec{Z}_{0},\vec{Z}_{1},\ldots,\vec{Z}_{T-1}$ denote i.i.d. random vectors in $\mathbb{R}^{N}$ following ${\cal N}(0,\frac{1}{N}{\bf I}_{N})$ distribution. For any unit $i$ and fixed $t\in\{0,\ldots,T-1\}$, we have
@@
where $R^{i}_{t}$ represents a random variable such that $\vec{R}_{t}:=\left(R^{1}_{t},\ldots,R^{N}_{t}\right)^{\top}=\sum_{i=0}^{t}\beta^{i}_{t}\vec{Z}_{i}$, with $\vec{\beta}_{t}=(\beta^{0}_{t},\ldots,\beta^{t}_{t},0,\ldots,0)^{\top}\in\mathbb{R}^{N}$ denoting a random vector that is correlated with $\vec{Z}_{0},\vec{Z}_{1},\ldots,\vec{Z}_{t}$ and satisfies ${\|\vec{\beta}_{t}\|}=1$. Furthermore,
@@
Equation (10) provides a decomposition of unit outcomes: the first term captures the interference effect, the second term reflects the unit-specific effect, the third term arises from uncertainties in the network structure, and the fourth term accounts for observation noise. We can then adjust the first term in Eq. (10) to reflect the available knowledge about the interference network.
Example 4.5
When it is known that no interference exists, Eq. (10) simplifies to $Y^{i}_{t+1}{}(\bm{W}_{t+1}{})=h_{t}{}\big{(}Y^{i}_{t}{}(\bm{W}_{t}{}),W^{i}_{t+1}{},{\vec{X}}^{i}\big{)}+\epsilon^{i}_{t}$, representing a general specification for the potential outcomes.
Example 4.6
Consider a scenario where the interference network is partially known and exhibits a clustered structure with clusters $C^{1},\ldots,C^{K}$ (e.g., the experimental population consists of individuals from different cities). In this context, Eq. (10) can be rewritten as follows:
$Y^{i}_{t+1}{}(\bm{W}_{t+1}{})=$ $\frac{1}{N}\sum_{k=1}^{K}\Bigg{[}\sum_{j\in C^{k},i\notin C^{k}}\left(\mu^{ij}+\mu_{t}^{ij}\right)g_{t}\left(Y^{j}_{t}{}(\bm{W}_{t}{}),W^{j}_{t+1}{},{\vec{X}}^{j}\right)$
$+\sum_{i,j\in C^{k}}\left(\mu^{ij}+\mu_{t}^{ij}\right)g_{t}\left(Y^{j}_{t}{}(\bm{W}_{t}{}),W^{j}_{t+1}{},{\vec{X}}^{j}\right)\Bigg{]}$
where we expect the magnitude of $\left(\mu^{ij}+\mu_{t}^{ij}\right)$ to be relatively larger whenever $i,j\in C^{k}$, indicating that units within the same cluster exhibit stronger ties.
The outcome decomposition rule in Eq. (10) leads to the following corollary, which characterizes the decomposition of the outcome sample mean for units belonging to subpopulation $\mathcal{S}$.

### v2

@@ 3.4.1 An Illustrative Example
Consider a special case of outcome specification (3) where functions $h_{t}$ and $g_{t}$ satisfy,
$g_{t}\big(Y^{i}_{t}{}(\bm{W}_{t}{}),W^{i}_{t+1},{\vec{X}}^{i}\big)=\delta_{t}+\theta_{t}Y^{i}_{t}{}(\bm{W}_{t}{})+\tau_{t}W^{i}_{t+1},\quad h_{t}\big(Y^{i}_{t}{}(\bm{W}_{t}{}),W^{i}_{t+1},{\vec{X}}^{i}\big)=\lambda_{t}Y^{i}_{t}{}(\bm{W}_{t}{})W^{i}_{t+1},$ (6)
where $\delta_{t},\theta_{t},\tau_{t},$ and $\lambda_{t}$ are unknown random variables independent of everything else.
Proposition 3.6
Suppose that Eq. (6) holds and for each unit $i$ and time $t$, and the empirical distributions of $(\mu^{i1},\ldots,\mu^{iN})$ and $(\mu^{i1}_{t},\ldots,\mu^{iN}_{t})$ converge weakly to the distributions of random variables M{M} and ${M}_{t}$, respectively, as $N\rightarrow\infty$. Additionally, assume the following holds for all $t$:
$\delta:=\mathbb{E}[({M}+{M}_{t})\delta_{t}],\quad\theta:=\mathbb{E}[({M}+{M}_{t})\theta_{t}],\quad\tau:=\mathbb{E}[({M}+{M}_{t})\tau_{t}],\quad\lambda:=\mathbb{E}[({M}+{M}_{t})\lambda_{t}].$ (7)
Then, the sample mean of outcomes $\sum_{i=1}^{N}Y^{i}_{t}{}(\bm{W}_{t}{})/N$ converges almost surely to a deterministic limit as $N\to\infty$. Denoting this limit by $\nu_{t}{}$, it satisfies the following dynamics:
@@
where $f(a,b)=\delta+\theta a+\tau b+\lambda ab$, that plays the role of the time-invariant function characterizing the evolution of outcomes over time. Moreover, if $\bm{w}{}$ and $\bm{w}{}^{\prime}$ share identical columns for treatment assignment at time $t=0$ and the set of marginal expectations of $\Pi_{T}$, $\{p_{0},\ldots,p_{T}\}$, contains at least two distinct values, Algorithm 1 provides consistent estimator for counterfactual evolutions.
Proposition 3.6 exemplifies our estimation strategy in a simple setting and presents the type of assumptions required to demonstrate consistency results. Specifically, according to Eq. (6) and Proposition 3.6, although we allow for varying interaction mechanisms across units and time, stability at the population level (see Eq. 7) ensures time-invariant evolution mechanisms as given in Eq. (8). We present a detailed analysis of a more general setting in Appendix 11.1.
@@
[…]
@@
3.5 Practical Limitations
The effectiveness of approaches like Algorithm 1 in identifying treatment effects across various settings has been shown empirically by Shirani and Bayati, (2024). However, two significant constraints require consideration. First, the estimation procedure in the second step of Algorithm 1 faces inherent sample size limitations that are bounded by the temporal horizon of the experiment. Second, as demonstrated by Bayati et al., (2024), certain settings require accounting for more complex structural relationships, necessitating a richer functional class for the outcome specification. Furthermore, estimation validity can be compromised by potential model misspecification (Karwa and Airoldi, 2018), highlighting the need for robust validation techniques.
These considerations raise two questions: First, how can we effectively expand the sample size to enhance estimation accuracy? Second, how can we determine the appropriate model specification for the underlying experiment, analogous to model selection procedures in supervised learning? The following two sections aim at addressing these questions.
@@ 4 Distribution-preserving Network Bootstrap
Our estimation strategy is grounded in the state evolution equation (4), as illustrated in Figure 3. In this section, we enhance our estimation capabilities by developing a resampling method that generates additional samples from the state evolution equation. The key observation underlying our approach is that the SE equation captures the dynamics of outcome distributions as described by Eq. (4). Therefore, different samples should provide different outcome distributions, but these distributions must evolve according to the same state evolution mappings $f_{t}$.
Our methodology batches experimental units into subpopulations, which we denote by $\mathcal{S}$. Each subpopulation $\mathcal{S}$ comprises a subset of units selected from $\{1,\ldots,N\}$, where membership is determined exclusively by the treatment allocation $\bm{W}{}$. Under the randomized treatment assumption, which specifies that $\bm{W}{}$ is random and independent of all other variables, each subpopulation $\mathcal{S}$ constitutes a random sample from the experimental population. This ensures sufficient similarity between distinct subpopulations to expect them to evolve under the same mappings. However, we must ensure that each particular configuration of $\mathcal{S}$ yields a unique observation of the SE equation. Consequently, we require that the empirical treatment distributions exhibit sufficient variation across subpopulations. This variation guarantees that different subpopulations provide genuinely distinct information about the underlying evolutionary dynamics.
We develop our distribution-preserving network bootstrap methodology through a sequence of theoretical results that build toward our main practical proposal. Specifically, we begin by extending the state evolution framework to characterize the dynamics of subpopulation-level quantities. We then introduce a finite-sample decomposition rule. This rule is central to our argument, as it mathematically grounds a fundamental bias-variance trade-off: it reveals how the unobserved mean and heterogeneity of the interference structure drive estimation bias and variance, respectively. This insight highlights the difficulty of determining an optimal estimation strategy ‘a priori’ and motivates the primary contribution of this section: a pragmatic framework that leverages DPNB as a tool to generate sufficient data for robust model selection via cross-validation. Finally, we show how this framework, developed for the general case of an unknown network, can be readily adapted to incorporate partial network knowledge for improved accuracy.
@@
The theoretical frameworks presented here extend to arbitrary subpopulations. While our analysis focuses specifically on treatment-allocation-based subpopulations in this section, the appendices provide complete theoretical statements applicable to generic subpopulations.
@@ 4.1 Subpopulation-specific State Evolution
We begin by extending the result of Eq. (4) to subpopulations of experimental units.
Theorem 4.2
@@
Consider a subpopulation $\mathcal{S}$ whose size grows to infinity as $N\to\infty$. Under Assumptions 3.3 and 3.3 and certain regularity conditions, for $t=0,1,\ldots,T-1$, we have
$H^{\mathcal{S}}_{0}{}$ $=h_{0}\big(Y^{\mathcal{S}}_{0}{},W^{\mathcal{S}}_{1},{\vec{X}}\big)+\epsilon_{0},$ (9)
$\tilde{\nu}_{1}{}$ $=\mathbb{E}\left[({M}+{M}_{0})g_{0}\big(Y_{0}{},W_{1},{\vec{X}}\big)\right],$
$(\tilde{\rho}_{1}{})^{2}$ $=(\sigma^{2}+\sigma_{0}^{2})\mathbb{E}\left[g_{0}\big(Y_{0}{},W_{1},{\vec{X}}\big)^{2}\right],$
$H^{\mathcal{S}}_{t}{}$ $=h_{t}\big(\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z_{t}+H^{\mathcal{S}}_{t-1}{},W^{\mathcal{S}}_{t+1},{\vec{X}}\big)+\epsilon_{t},$
$\tilde{\nu}_{t+1}{}$ $=\mathbb{E}\left[({M}+{M}_{t})g_{t}\big(\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z_{t}+H_{t-1}{},W_{t+1},{\vec{X}}\big)\right],$
$(\tilde{\rho}_{t+1}{})^{2}$ $=(\sigma^{2}+\sigma_{t}^{2})\mathbb{E}\left[g_{t}\big(\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z_{t}+H_{t-1}{},W_{t+1},{\vec{X}}\big)^{2}\right],$
where
  • $Y_{0}{}$ and $Y^{\mathcal{S}}_{0}{}$ represent the weak limits of the population and subpopulation initial outcomes;
  • $W_{t}$ and $W^{\mathcal{S}}_{t}$ are the weak limits of the population and subpopulation treatment assignments;
  • $\vec{X}$} represents the weak limit of the population covariates;
  • M{M} and ${M}_{t}$ represent the weak limits of $(\mu^{i1},\ldots,\mu^{iN})$ and $(\mu^{i1}_{t},\ldots,\mu^{iN}_{t})$, respectively;
  • $\epsilon_{t}$ is the weak limit of noise terms $(\epsilon^{1}_{t},\ldots,\epsilon^{N}_{t})$;
  • $Z_{t}$ follows a standard Gaussian distribution;
  • $H_{t}{}$ denotes the population-level counterpart of $H^{\mathcal{S}}_{t}{}$, when $\mathcal{S}$ is equal to the entire population.
Furthermore, if we denote by $Y^{\mathcal{S}}_{t}{}$ the weak limit of outcomes for units belonging to $\mathcal{S}$ as $N\rightarrow\infty$, then $Y^{\mathcal{S}}_{t}{}$ follows the same distribution as $\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z_{t}+H^{\mathcal{S}}_{t-1}{}$. Similarly, denoting by $Y_{t}{}$ the weak limit of outcomes at the population level, $Y_{t}{}$ follows the same distribution as $\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z_{t}+H_{t-1}{}$.
We can express the state evolution equation (9) compactly using mappings $\varphi_{t},\;t=0,\ldots,T-1$:
$Y^{\mathcal{S}}_{t+1}{}\stackrel{{\scriptstyle d}}{{=}}\varphi_{t}\left(Y_{t}{},W_{t+1}{},Y^{\mathcal{S}}_{t}{},W^{\mathcal{S}}_{t+1}{}\right),$ (10)
@@
where $\stackrel{{\scriptstyle d}}{{=}}$ refers to equality in distribution and each mapping $\varphi_{t}$ involves the functions $g_{t}$ and $h_{t}$, the random objects M{M}, ${M}_{t}$, $Z_{t+1}$, $\vec{X}$} and $\epsilon_{t}$, as well as the constants $\sigma$ and $\sigma_{t}$. Then, by Eq. (10), the outcomes of units within $\mathcal{S}$ evolve through a dynamic interplay between population-level and subpopulation-specific outcomes and treatments. The formal and more general statement of Theorem 4.2 is presented in Theorem 10.9 of Appendix 10.4.
@@
To demonstrate the efficacy of DPNB, consider two subpopulations $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ with distinct treatment allocations ($W^{\mathcal{S}_{1}}_{t+1}{}\neq W^{\mathcal{S}_{2}}_{t+1}{}$ in Eq. (10)). Then, two sequences $Y^{\mathcal{S}_{1}}_{0}{},\ldots,Y^{\mathcal{S}_{1}}_{T}{}$ and $Y^{\mathcal{S}_{2}}_{0}{},\ldots,Y^{\mathcal{S}_{2}}_{T}{}$ yield different samples of the state evolution described by (10). Notably, for each time step $t$, the mapping $\varphi_{t}$ shares fundamental structural properties with $f_{t}$, with the experimental state evolution equation in (4) emerging as a special case of Eq. (10) when applied to the entire population. This implies that by strategically selecting subpopulations with varying treatment histories, we obtain multiple state evolution trajectories, facilitating the estimation of $\varphi_{t}$ and, consequently, $f_{t}$.
@@
Consider two distinct subpopulations based on treatment assignment $W_{t+1}{}$: treated units ($\mathcal{S}={\cal T}$) and control units ($\mathcal{S}={\cal C}$). From Eq. (10), we have:
$Y^{{\cal T}}_{t+1}{}\stackrel{{\scriptstyle d}}{{=}}\varphi_{t}\left(Y_{t}{},W_{t+1}{},Y^{{\cal T}}_{t}{},1\right),\quad\quad Y^{{\cal C}}_{t+1}{}\stackrel{{\scriptstyle d}}{{=}}\varphi_{t}\left(Y_{t}{},W_{t+1}{},Y^{{\cal C}}_{t}{},0\right).$ (11)
When the direct treatment effect is non-zero, the outcome distributions for the treated group $Y^{{\cal T}}_{t}{}$ and the control group $Y^{{\cal C}}_{t}{}$ are distinct. Nevertheless, by (11), both distributions evolve according to the same mappings $\varphi_{t}$, thereby providing two different observations from the state evolution.
Example 4.3 demonstrates how DPNB enables creating multiple observations of the state evolution equation with different treatment intensity levels. Now, we show that DPNB indeed generates representative samples to effectively support robust estimation of state evolution mappings. To illustrate, we must first introduce a theoretical result, a finite-sample analysis that yields a structural decomposition of the outcome specification. This decomposition disentangles the network (or indirect) treatment effect from the direct effect and identifies an endogenous noise term whose magnitude reflects the heterogeneity of the underlying network structure. Then, we discuss how DPNB provides a way to reduce the impact of this endogenous noise through averaging, while simultaneously generating sufficient representative samples for reliable estimation.
4.2 Outcomes Decomposition and Its Implications
@@
Suppose Assumption 3.3 holds. For any unit $i$ and fixed $t\in\{0,\ldots,T-1\}$, we have
@@
where $R^{i}_{t}$ is a random variable such that $\vec{R}_{t}:=\left(R^{1}_{t},\ldots,R^{N}_{t}\right)^{\top}=\sum_{i=0}^{t}\beta^{i}_{t}\vec{Z}_{i}$, with $\vec{Z}_{0},\vec{Z}_{1},\ldots,\vec{Z}_{T-1}\in\mathbb{R}^{N}$ are i.i.d. random vectors following ${\cal N}(0,\frac{1}{N}{\bf I}_{N})$ distribution and $\vec{\beta}_{t}=(\beta^{0}_{t},\ldots,\beta^{t}_{t},0,\ldots,0)^{\top}\in\mathbb{R}^{N}$ denotes a random vector that is correlated with $\vec{Z}_{0},\vec{Z}_{1},\ldots,\vec{Z}_{t}$ and satisfies ${\|\vec{\beta}_{t}\|}=1$. Furthermore,
@@
We now discuss several implications of Theorem 4.4. For expositional clarity, we assume a time-invariant network structure and set $\mu^{ij}_{t}=0$ and $\sigma_{t}=0$ for all $t$, while supposing $\mu^{ij}\equiv\mu$ for all $i,j$. Then, the only heterogeneity in the network structure arises from i.i.d. Gaussian random variables with variance $\sigma^{2}/N$, while $\mu$ reflects the average intensity of interactions (see Assumptions 3.3). Omitting the explicit outcome–treatment dependence in the notation, Eq. (12) reduces to
$Y^{i}_{t+1}{}=\underbrace{\mu\,\bar{G}_{t}}_{\text{global interference}}\;+\;\underbrace{h_{t}\big(Y^{i}_{t}{},W^{i}_{t+1}{},{\vec{X}}^{i}\big)}_{\text{unit/direct term}}\;+\;\underbrace{\sigma\,{\left\|g_{t}\!\left(\vec{Y}_{t}{},\vec{W}_{t+1}{},\bm{X}\right)\right\|}R^{i}_{t}}_{\text{network heterogeneity noise}}\;+\;\epsilon^{i}_{t},$ (13)
where $\bar{G}_{t}:=\frac{1}{N}\sum_{j=1}^{N}g_{t}\!\big(Y^{j}_{t}{},W^{j}_{t+1}{},{\vec{X}}^{j}\big)$. By Theorem 4.4, we can approximately view $R^{i}_{t}$ as an additive, mean-zero perturbation with variance of order $1/N$. Consequently, the third term in (13) behaves like additional noise with variance
$\frac{\sigma^{2}}{N}\,{\left\|g_{t}\!\left(\vec{Y}_{t}{},\vec{W}_{t+1}{},\bm{X}\right)\right\|}^{2}.$ (14)
Note that ${\left\|g_{t}\!\left(\vec{Y}_{t}{},\vec{W}_{t+1}{},\bm{X}\right)\right\|}^{2}$ also scales with $N$, implying that the term in Eq. (14) does not vanish as $N\to\infty$. This persistence of noise has direct implications for standard estimators such as difference-in-means (DM), which is defined at time $t+1$ as follows:
$\widehat{\tau}^{\,\mathrm{DM}}_{t+1}=\frac{1}{n_{\cal T}}\sum_{i\in{\cal T}}Y^{i}_{t+1}{}-\frac{1}{n_{\cal C}}\sum_{i\in{\cal C}}Y^{i}_{t+1}{},$
with $\cal T$} and $\cal C$} representing the treatment and control groups, respectively. In Eq. (13), the global interference term is equal to $\mu\,\bar{G}_{t}$ for every $i$ and cancels in the difference. Then, $\widehat{\tau}^{\,\mathrm{DM}}_{t+1}$ behaves “as if” there were no interference but with inflated noise. Using (14) and the approximation $R^{i}_{t}\stackrel{{\scriptstyle\mathrm{approx}}}{{\sim}}{\cal N}(0,1/N)$, the interference-induced variance inflation of $\widehat{\tau}^{\,\mathrm{DM}}_{t}$ is approximately
$\frac{\sigma^{2}}{N}\,{\left\|g_{t}\!\left(\vec{Y}_{t}{},\vec{W}_{t+1}{},\bm{X}\right)\right\|}^{2}\left(\frac{1}{n_{\cal T}}+\frac{1}{n_{\cal C}}\right),$ (15)
in addition to the variance contributed by observation noise $\epsilon^{i}_{t}$. That means the heterogeneity of the (unknown) interference, captured by $\sigma$, acts as extra noise even when the estimator ignores interference. However, the price of “ignoring” $\mu\,\bar{G}_{t}$ is that $\widehat{\tau}^{\,\mathrm{DM}}_{t+1}$ incurs a bias in estimating any counterfactual $\text{CFE}_{t+1}{(\bm{w}{}^{\prime})}$, depending on the magnitude of $\mu\,\bar{G}_{t}$ at treatment level $\bm{w}{}^{\prime}$.
These observations lead to the following takeaways. First, $\mu$ primarily drives bias, whereas $\sigma$ primarily drives variance through network heterogeneity noise. Second, structural knowledge of the average interaction term captured by $\mu$ enables us to estimate and subtract the first term in (12) and reducing bias. However, this does not reduce the variance contribution in (15), as it originates from $\sigma$ and the near-Gaussian perturbations $\vec{R}_{t}$. We illustrate these phenomena through a simple simulation that varies $\mu$ and $\sigma$, and then discuss how to address the bias component.
We consider a simple outcome specification where $h_{t}\big(Y^{i}_{t}{}(\bm{W}_{t}{}),W^{i}_{t+1},{\vec{X}}^{i}\big)=1-1.2W^{i}_{t+1}$ and $g_{t}\big(Y^{i}_{t}{}(\bm{W}_{t}{}),W^{i}_{t+1},{\vec{X}}^{i}\big)=W^{i}_{t+1}$. The experiment runs for eight time periods with a population of $N=500$ units. The treatment assignment for each unit follows a Bernoulli distribution, with probability $0.25$ for the first four periods and $0.75$ for the remaining four periods. Finally, the standard deviation of the observation noise is set to $\sigma_{\epsilon}=0.1$.
We perform two sets of simulations on the effect of interference matrix parameters $\mu$ and $\sigma$ and display the results in Figure 4. In the first set (left panel), we fix $\mu=0.04$ and vary $\sigma\in\{0.1,,0.2,,0.4,,0.8,,1.6\}$. In the second set (right panel), we fix $\sigma=0.5$ and vary $\mu\in\{0.01,,0.02,,0.04,,0.08,,0.16,,0.32\}$. For each configuration, we compute the ground‑truth TTE and the DM estimate at the final time period. To quantify uncertainty and decompose error, we use a nested bootstrap: we simulate $100$ independent “worlds” (fresh interference matrix $A$ and initial outcomes $Y_{0}$) and $200$ resamples of treatment allocations $\bm{W}{}$ per world; we then compute $\pm 1$ SE bands for Mean Squared Error (MSE), variance, and squared bias of TTE estimation using $400$ nested bootstrap resamples that re‑sample worlds and runs.
Figure 4: Difference‑in‑Means (DM) performance under interference. Left: $\sigma$ sweep at fixed $\mu=0.04$. Right: $\mu$ sweep at fixed $\sigma=0.5$. Curves show MSE, variance, and squared bias with $\pm 1$ SE bands from a nested bootstrap. Axes are log–log.
The simulation results in Figure 4 empirically validate our theoretical findings. The left panel shows that when $\mu$ is fixed and $\sigma$ increases, the variance of the DM estimator grows, while its squared bias remains constant. This demonstrates that network heterogeneity, parameterized by $\sigma$, inflates the variance of simple estimators. In contrast, the right panel shows that when $\sigma$ is fixed and $\mu$ increases, the squared bias grows, dominating the MSE.
4.3 Navigating Estimation Trade-offs with DPNB
The preceding analysis highlights the limitations of standard estimators such as DM. While variance inflation from network heterogeneity ($\sigma$) reduces statistical power, the bias from mean interference ($\mu$) poses a more fundamental problem that can lead to incorrect causal conclusions. Strategies like Algorithm 1 address this bias by modeling the unknown interference structure. However, such population-level approaches are limited by the sample size $T$. To utilize the full statistical power across all $N\times T$ observations, we must apply DPNB and develop a more sophisticated estimation framework. Below, our subpopulation decomposition rule provides the precise theoretical tool to understand the statistical trade-offs inherent in this task. It characterizes how averaging over subpopulations of different sizes affects the endogenous noise introduced by network heterogeneity.